If y1 = 2x + 2and y2 = −x^2/2 are the solutions of the equation y = x y′+(y')^2/ 2then are the constant multiples c1y1and c2y2 , where c1, c2 are arbitrary, also the solutions of the given de? is the sum y1+ y2 a solution? Justify your answer.
1
Expert's answer
2019-04-22T10:23:20-0400

Let


"y_1=2x+2, y_2=-{x^2 \over 2}"

Consider


"y^*=c_1y_1+c_2y_2"

"y^*=c_1(2x+2)+c_2(-{x^2 \over 2})=(-{c_2 \over 2})x^2+(2c_1)x+2c_1"

where c1, c2 are arbitrary constants.

Then


"(y^*)'=c_1(y_1)'+c_2(y_2)'=c_1(2x+2)'+c_2(-{x^2 \over 2})'=2c_1-xc_2"

Consider


"x(y^*)'+{((y^*)')^2 \over 2}"

"x(2c_1-xc_2)+{(2c_1-xc_2)^2 \over 2}=2xc_1-x^2c_2+2c_1^2-2xc_1c_2+{1 \over 2}x^2c_2^2="

"=({1 \over 2}c_2^2-c_2)x^2+(2c_1-2c_1c_2)x+2c_1^2"

The necessary conditions are


"{1 \over 2}c_2^2-c_2=-{c_2^2 \over 2}""2c_1-2c_1c_2=2c_1""2c_1^2=2c_1"

"c_2(c_2-1)=0""c_1c_2=0""c_1(c_1-1)=0"

If "c_1=0," then "c_2=0" or "c_2=1".

If "c_2=0," then "c_1=0" or "c_1=1."

We have three pairs for "(c_1, c_2)"


"(0, 0), (0, 1), (1, 0)"

The sum "y_1+y_2" is not the solution of the differential equation.